Calculate the De Broglie wavelength of an electron with an energy of 487eV

Give your answers in nm to 3 decimal place
Answer #1

(amk De-Broglie wavelength (1) h m= mass of electron = 9.1 x 10-31 Kg K= Kinetic energy = 487 ev n = planers constant 6.6266.626 X10-34 J.S so, A= 2x 9-1X10-31 kg x 487% 1.6X10-19 J ::de 6.626 X 10-34 m H1008 x 10-23 id=5.543 X 10-11 m :1-0.055 X10